OpenStudy (anonymous):

If the ratio of similarity between two similar polygons is 4:3.2, this ratio converts to 1:8. 8:1. 4:5. 5:4.

OpenStudy (raden):

4 : 3.2 = 4 : 32/10 = 4 : 16/5 change into multiplication = 4 x 5/16 = 20 / 16 = 5 : 4

OpenStudy (anonymous):

I still dont understand how to get the answer and I have more problems that are similar. Can You please explain it

OpenStudy (raden):

just convert 3.2 as a fraction, it is 32/10

OpenStudy (raden):

and it simplied be 32/10 = 16/5

OpenStudy (raden):